For this problem, carry at least four digits after the decimal in your calculations. Answers may vary slightly due to rounding.

In a random sample of 502 judges, it was found that 278 were introverts.

(b) Find a 99% confidence interval for p. (Round your answers to two decimal places.)

lower limit    
upper limit    
Answer #1

Solution :

Given that,

n = 502

x = 278

Point estimate = sample proportion = \hat p = x / n = 278/502=0.554

1 - \hat p   = 1- 0.554 =0.446

At 99% confidence level the z is ,

\alpha  = 1 - 99% = 1 - 0.99 = 0.01

\alpha / 2 = 0.01 / 2 = 0.005

Z\alpha/2 = Z0.005 = 2.576 ( Using z table )

  Margin of error = E = Z\alpha / 2    * (\sqrt(((\hat p * (1 - \hat p )) / n)

= 2.576* (\sqrt((0.554*0.446) /502 )

E = 0.06

A 99% confidence interval for population proportion p is ,

\hat p - E < p < \hat p + E

0.554 - 0.06< p < 0.554+0.06

0.49< p < 0.61

The 99% confidence interval for the population proportion p is : lower limit=0.49,upper limit= 0.61